solve the equation with the quadratic formula:
9m^2=-2

Answer by RAY100(1637)

9m^2 =-2
.
9m^2 +0m +2=0
.
a=9,,,b=0,,,c=2
.
m=[ -(0)+/- sqrt{(0)^2 -4 (9)(2)} ] / 2(9)
.
m= +/- sqrt (-72) /18
.
m= +/- .471 i
.
check
9m^2 = -2
m^2 = -2/9
m=+/- sqrt ( -2/9) = +/- .471 i,,,,,,ok